# S3c-Math: Spontaneous Step-level Self-correction Makes Large Language Models Better Mathematical Reasoners

### v1

@@ Introduction
Reasoning is one of the essential foundational abilities of large language models (LLMs), showing the capacity of LLMs to tackle complex real-world problems. Nowadays, researchers are increasingly focusing on the performance of LLMs in specific reasoning tasks like mathematics, code, logic, common-sense, etc. (Sun et al. 2024). Mathematics is one of the significant branches of reasoning ability of LLMs, and solving a mathematical problem can demonstrate an LLM’s ability to decompose, reason, and summarize complex problems. In recent work, the chain-of-thought (CoT) inference (Wei et al. 2023) has been proven to be a method that can significantly enhance an LLM’s ability to solve reasoning problems (Suzgun et al. 2022; Wang and Zhou 2024). CoT induces the model to explicitly output the reasoning process, gradually deriving a series of intermediate processes or sub-goals, thereby enabling the model to correctly answer reasoning questions. However, during the step-by-step reasoning process, an LLM may still generate errors (Chen et al. 2024; Wang et al. 2024). These errors can be propagated to subsequent reasoning processes, leading to incorrect outputs from the LLMs.
@@
Next, we will use the trained model to evaluate the correctness of the generated candidates $(\hat{x}_{i+1})$. Specifically, for each step $\hat{x}_{i+1}$ in the candidate set, we concatenate the correct preceding steps $x_{1},\ldots,x_{i}$ as context before the step $\hat{x}_{i+1}$, and ask the model $M_{s}$ to continue generating until the model actively ends the generation (i.e., encounters the ¡end-of-sequence¿ token). We will determine whether the step is correct based on whether the answer generated by model $M_{s}$ matches the standard answer. In order to make more accurate judgments and reduce the cases where correct steps are mistakenly considered as wrong, we use $pass@k$ to evaluate a step. Specifically, we sample multiple model outputs with a high temperature of 1.0 for 16 times, and only when none of the contents generated by the model match the correct answer do we consider the step to be wrong. Whether $\hat{x}_{i+1}$ is correct can be represented as:
@@
Based on the correct steps in the existing data and the erroneous steps we sampled in the previous section, we inserted the erroneous steps and the flags used to trigger self-correction into the correct steps, thus forming the direct self-correction data. However, another significant challenge of self-correction is how the LLMs can produce more accurate content when they know that errors already exist. To enable the LLMs to better learn to correct the identified mistakes, we annotated the acquired step-level self-correction samples in two ways, reflection and improvement, using Meta-Llama-3-70B-Instruct (Meta AI 2024). Reflection involves the model analyzing where errors occurred in the existing answers, while improvement entails generating ways to improve based on the existing output and reflection. When generating these two annotations, we provide the model with a prompt, as well as a question, previous step, wrong step, and correct step, and ask the model to make annotations based on this information. The prompt we used is shown in Appendix A. We present a data example and its components in Figure 3.
@@
Meta-Llama-3-8B MetaMathQA 81.12 86.20 30.58 39.92 81.40 86.60 18.27 28.21
@@
Meta-Llama-3-70B MetaMathQA 88.55 91.74 45.70 55.40 87.40 91.00 29.08 40.30
@@
DeepSeek-Math-Base-7B MetaMathQA 79.30 85.90 38.22 46.86 80.60 86.40 25.98 42.95
@@
Qwen2-Math-7B MetaMathQA 84.08 89.08 51.32 60.64 85.60 90.40 42.44 52.02
@@
Table 1: Our main experimental results (%) on four mathematical reasoning tasks (GSM8K, MATH, SVAMP and Mathematics) under $Pass@1$ and $Majority@32$ settings. The term ”w/o. R&I” represents that the self-correction part of the SFT samples do not include reflection and improvements, while ”w/o. I” indicates that the self-correction part of the SFT samples do not include improvements. We abbreviate $Pass@1$ as $P@1$ and $Majority@32$ as $M@32$. The highest accuracy is indicated in bold, while the second highest is underlined.
@@
We test the performance of our models on the evaluation sets of GSM8K (Cobbe et al. 2021), MATH (Hendrycks et al. 2021), SVAMP (Patel, Bhattamishra, and Goyal 2021), and Mathematics (Saxton et al. 2019). For each sample in the test set, we only perform inference once. Each test sample in these datasets contains a question and a golden answer. We extract the final answer from the model’s output and matched it with the golden answer to ultimately determine whether the sample was answered correctly. The evaluation uses the vllm framework  (Kwon et al. 2023) for inference, and we evaluate the model’s single-inference accuracy, i.e., the $pass@1$ metric, using greedy decoding. Self-consistency (Wang et al. 2023) is considered a stable way to evaluate the model’s sampling method. To more accurately assess the improvements brought about by our strategy, we also test the accuracy under the $maj@32$ setting, setting the temperature to 0.7 during sampling decoding.
@@
As shown in Table 2, the experiment demonstrates that altering the distribution of the query does not enhance the model’s performance, further validating the effectiveness of our method.
@@
To more clearly demonstrate the spontaneous step-level self-correction ability generated by our method in LLMs, we provide an evaluation output of a model trained on the DeepSeek-Math base. We present the results of greedy decoding produced by training on MetaMathQA and S3c-MathQA respectively in Table 3. We can see that after training on S3c-MathQA, the model spontaneously realizes its mistake in the second step where it incorrectly assumed ‘the rest of its life‘ to represent 365 days, and proceeds to recalculate, arriving at the final result. This case can also show that our method does not affect the original data to the greatest extent, as the models trained on these two sets of data produce almost identical greedy outputs on the same case in the first two steps.
@@ Conclusion
Appendix A. Prompts
In our paper, we employ Meta-Llama-3-70B-Instruct (Meta AI 2024) for generating reflection and improvement contents based on self-correction. Figure 4 showcases the prompts we utilized.

### v2

@@ Introduction
Reasoning is one of the essential foundational abilities of large language models (LLMs), showing the capacity of LLMs to tackle complex real-world problems. Nowadays, researchers are increasingly focusing on the performance of LLMs in specific reasoning tasks like mathematics, code, logic, common-sense, etc. (Sun et al. 2024). Mathematics is one of the significant branches of reasoning ability of LLMs, and solving a mathematical problem can demonstrate an LLM’s ability to decompose, reason, and summarize complex problems. In recent works, the chain-of-thought (CoT) inference (Wei et al. 2023) has been proven to be a method that can significantly enhance an LLM’s ability to solve reasoning problems (Suzgun et al. 2022; Wang and Zhou 2024). CoT induces the model to explicitly output the reasoning process, gradually deriving a series of intermediate processes or sub-goals, thereby enabling the model to correctly answer reasoning questions. However, during the step-by-step reasoning process, an LLM may still generate errors (Chen et al. 2024; Wang et al. 2024). These errors can be propagated to subsequent reasoning processes, leading to incorrect outputs from the LLMs.
@@
Next, we will use the trained model to evaluate the correctness of the generated candidates $(\hat{x}_{i+1})$. Specifically, for each step $\hat{x}_{i+1}$ in the candidate set, we concatenate the correct preceding steps $x_{1},\ldots,x_{i}$ as context before the step $\hat{x}_{i+1}$, and ask the model $M_{s}$ to continue generating until the model actively ends the generation (i.e., encounters the [end-of-sequence] token). We will determine whether the step is correct based on whether the answer generated by model $M_{s}$ matches the standard answer. In order to make more accurate judgments and reduce the cases where correct steps are mistakenly considered as wrong, we use $pass@k$ to evaluate a step. Specifically, we sample multiple model outputs with a high temperature of 1.0 for 16 times, and only when none of the contents generated by the model match the correct answer do we consider the step to be wrong. Whether $\hat{x}_{i+1}$ is correct can be represented as:
@@
Based on the correct steps in the existing data and the erroneous steps we sampled in the previous section, we inserted the erroneous steps and the flags used to trigger self-correction into the correct steps, thus forming the direct self-correction data. However, another significant challenge of self-correction is how the LLMs can produce more accurate content when they know that errors already exist. To enable the LLMs to better learn to correct the identified mistakes, we annotated the acquired step-level self-correction samples in two ways, reflection and improvement, using Meta-Llama-3-70B-Instruct (Meta AI 2024). Reflection involves the model analyzing where errors occurred in the existing answers, while improvement entails generating ways to improve based on the existing output and reflection. When generating these two annotations, we provide the model with a prompt, as well as a question, previous steps, wrong step, and correct step, and ask the model to make annotations based on this information. We present a data example and its components in Figure 3.
@@
Meta-Llama-3-8B MetaMath 81.12 86.20 30.58 39.92 81.40 86.60 18.27 28.21
@@
Meta-Llama-3-70B MetaMath 88.55 91.74 45.70 55.40 87.40 91.00 29.08 40.30
@@
Deepseek-math-base MetaMath 79.30 85.90 38.22 46.86 80.60 86.40 25.98 42.95
@@
Qwen2-Math-7B MetaMath 84.08 89.08 51.32 60.64 85.60 90.40 42.44 52.02
@@
Table 1: Our main experimental results (%) on four mathematical reasoning tasks (GSM8K, MATH, SVAMP and Mathematics) under $Pass@1$ and $Majority@32$ settings. The term ”w/o. R&I” represents that the self-correction part of the SFT samples do not include reflection and improvements. We abbreviate $Pass@1$ as $P@1$ and $Majority@32$ as $M@32$. The highest accuracy is indicated in bold, while the second highest is underlined.
@@
We test the performance of our models on the evaluation sets of GSM8K (Cobbe et al. 2021), MATH (Hendrycks et al. 2021), SVAMP (Patel, Bhattamishra, and Goyal 2021), and Mathematics (Saxton et al. 2019). Each test sample in these datasets contains a question and a golden answer. We extract the final answer from the model’s output and matched it with the golden answer to ultimately determine whether the sample was answered correctly. The evaluation uses the vllm framework  (Kwon et al. 2023) for inference, and we evaluate the model’s single-inference accuracy, i.e., the $pass@1$ metric, using greedy decoding. Self-consistency (Wang et al. 2023) is considered a stable way to evaluate the model’s sampling method. To more accurately assess the improvements brought about by our strategy, we also test the accuracy under the $maj@32$ setting, setting the temperature to 0.7 during sampling decoding.
@@
As shown in Table 2, altering the distribution of the query does not enhance the model’s performance, further validating the effectiveness of our method.
@@
To more clearly demonstrate the spontaneous step-level self-correction ability generated by our method in LLMs, we provide an evaluation output o25 f a model trained on the DeepSeek-Math base. We present the results of greedy decoding produced by training on MetaMathQA and S3c-MathQA respectively in Table 3. We can see that after training on S3c-MathQA, the model spontaneously realizes its mistake in the second step where it incorrectly assumed ‘the rest of its life‘ to represent 365 days, and proceeds to recalculate, arriving at the final result. This case can also show that our method does not affect the original data to the greatest extent, as the models trained on these two sets of data produce almost identical greedy outputs on the same case in the first two steps.
@@ Conclusion
Acknowledgments
This work is supported by National Natural Science Foundation of China (No. 62436007), the Key Research and Development Program of Zhejiang Province, China (No. 2023C01152) and ZJU Kunpeng&Ascend Center of Excellence.